OpenStudy (anonymous):

Prove 2SinAsinB =cos(A-B)-cos(A+B) FROM LHS <

OpenStudy (anonymous):

what u mean by LHS and RHS ? u can expand cos (A-B) and cos (a+b ) and sub their common terms . u will find the answer

OpenStudy (anonymous):

Prove from Left hand side saying what formula you have used? I know right hand side is addition formula.Prove starting from left side of eqn

OpenStudy (anonymous):

Yes we can do that just do opposite steps...

OpenStudy (anonymous):

Show what formula you have used and how you manipulate the eqn from LHS?

OpenStudy (anonymous):

Can you write \(2sinAsinB = sinAsinB + sinAsinB \) ???

OpenStudy (anonymous):

dear sinAsinB = 1/2(cos ( A-B)-cos(A+B)) just this

OpenStudy (anonymous):

LHS = \(sinAsinB + sinAsinB\) Or: \[\large = cosAcosB +sinAsinB - (cosAcosB - sinAsinB)\] \[\large = \cos(A-B) -\cos(A+B) \implies RHS\]

OpenStudy (anonymous):

tnx

OpenStudy (anonymous):

Can you guys do it step by step stating the formula you have used like "sin addition formula" instead of just showing me the working derived from right side?

OpenStudy (anonymous):

I am not getting what you said @vadevalor

OpenStudy (anonymous):

Can you state what formula you used to explain what you did for each step?

OpenStudy (anonymous):

Firstly I write the LHS as: \[\large sinAsinB + sinAsinB\] Got??

OpenStudy (anonymous):

Then add or subtract by \(cosAcosB\): \[\large cosAcosB + sinAsinB - cosAcosB + sinAsinB\] Group them now: \[\large (cosAcosB+sinAsinB)−(cosAcosB-sinAsinB)\] Now use the formula: \[\large = \cos(A-B) - \cos(A + B)\]

OpenStudy (anonymous):

The formula we will be using in this are: \[\large \color{green}{\cos(A + B) = cosA cosB - sinAsinB}\] \[\large \color{green}{\cos(A - B) = cosA cosB + sinAsinB}\]

OpenStudy (anonymous):

How do you transit from LHS to 1st step?

OpenStudy (anonymous):

2x you can write it as x + x or not???

OpenStudy (anonymous):

Okay got it thanks...sorry

OpenStudy (anonymous):

Welcome dear...